OpenStudy (anonymous):

Find the reference angle for the given angle θ. 5.45

OpenStudy (jdoe0001):

5.45 radians?

OpenStudy (anonymous):

I don't know all it says on the problem is 5.45. Should I assume it is radians?

OpenStudy (jdoe0001):

well.... could also be 5.45 degrees I gather

OpenStudy (anonymous):

the other problems all had a degree symbol next to them. I'm just gunna work it like it is radiams. Do you know how to work it if it is (8pi)/9?

OpenStudy (jdoe0001):

a reference angle, is just an angle that you could make with the horizontal line for example 

OpenStudy (jdoe0001):

5.45 radians a circle has an angle around of \(2\pi\) that is in radians in \(\pi\) terms \(\bf 2\pi\to 2\cdot 3.1416 \approx 6.2832 \\ \quad \\ 6.2832 - 5.45= 0.8332\ radians\)

OpenStudy (jdoe0001):

so.... the issue is, where is 5.45 radians at and I think is in the 3rd quadrant

OpenStudy (jdoe0001):

if you want to say... will be easier to find it with degrees if you want to convert it to degrees, all you have to do is \(\bf degrees=\cfrac{5.45\cdot 180}{\pi}\approx 312.26^o\) so where is 312.26 degrees? well